# positioning

### v1

@@ IV-B Model Training
We train both models on the classification task of predicting whether a viewpoint will result in localization errors smaller than 0.1m and 1 deg. The main pre-processing steps consist of scaling all of the input features $\mathbf{F}$ to values between 0 and 1 and balancing the data such that it contains roughly equal numbers of positive and negative samples. Split of the training and validation sets can be found in Figure 4. For each scene, 100 waypoints are randomly generated. At each waypoint, 50 viewpoints are sampled, allowing full rotation in the yaw axis, while limiting the pitch to be within the range of [-10deg, 45deg]. This creates a training+validation set of 25’000 images and a test set of 20’000 images.
[…]
Figure 4: Overview of the chosen Matterport environments. The number in brackets following the designation corresponds to the index in the Matterport dataset.
@@ IV-C Evaluation Strategy
As introduced in III, we evaluate three baseline methods $\uppi_{\textrm{FIM}},\uppi_{\textrm{max}},\uppi_{\textrm{angle}}$. We supplement them with two extra naive approaches: $\uppi_{\textrm{forward}}$, which simply points the camera towards the next waypoint along the trajectory, and $\uppi_{\textrm{random}}$, which randomly selects the viewpoints. Similar to the training process, for each evaluation environment, random waypoints and samples of viewpoints are generated. All sampled viewpoints are then evaluated by each viewpoint-choosing strategy, saving the localization error of each viewpoint chosen. As an upper bound to the achievable localization accuracy for a given position, the smallest error of the hloc localization for each is also saved and considered the ”optimal” viewpoint-choosing strategy.
@@ IV-D Real-world Deployment and Test
Setup To further validate the results obtained from the simulation experiments, the best-performing methods were implemented as a ROS-compatible planning module and deployed on a quadruped robot with a robotic arm that contains a calibrated color camera in the end effector, allowing different viewpoints to be viewed for a given body position. We use a HoloLens 2 to build a map of an indoor environment, as shown in Figure 5.
Data Collection Unlike the simulation, we no longer have the ground-truth poses in a real-world deployment. To evaluate the viewpoint-choosing strategies, we require two types of positional information: Each viewpoint planner needs to get an initial estimate $\hat{\bm{p}}$ that is used in Equation 3 to select a viewing direction at the current location. Given that viewing direction, the robot moves the arm and takes an image, and localizes this with respect to the existing map. We then need a ground-truth pose to evaluate this result against.
To measure accurate poses in the map created from the HoloLens recording, we use a combination of hloc and AprilTag fiducial markers [21]. Once a map of the environment has been created, an AprilTag is affixed to a static location. We capture several high-quality images of the marker, including its surroundings, using a calibrated DSLR camera. The AprilTag marker can be accurately localized within the captured images, and by ensuring that the images contain enough of the surroundings, they can also be localized in the map using hloc. This two-step process provides an estimated location of the AprilTag marker within the map for each image. An average of the estimated locations is taken in order to minimize the effect of errors in the tag or image localizations. This process gives us accurate poses of fiducial markers with respect to the captured landmark map.
@@
A similar approach can then be used to obtain a best-possible estimate of the robot body pose at any position in the map by placing a marker on a stable tripod that can be detected in both the robot’s end-effector camera and in a collection of calibrated DSLR images. We use the body pose of the robot estimated in that way, together with the inverse kinematics of the arm, as ground truth that we compare the hloc results to.
@@ V Results
Figure 6: Cumulative distribution of position errors of the evaluation points in the testing environments using SuperPoint(Left) and SIFT(Right) local features. Methods or versions of methods that rely on the environment mesh in order to determine landmark occlusion are plotted using normal-⋅normal-$\cdot$-$\cdot$-⋅ - ⋅ - lines.
As introduced in the last section, we evaluate our approaches and the baseline methods in both our simulation pipeline and the real world. Table I shows the quantitative results of the simulation. To demonstrate the generalization ability of our approaches across different feature descriptors, we evaluate all methods using both SuperPoint feature [9] and SIFT feature [19], whereas both of our scoring models are only trained with the SuperPoint feature. The result in the table shows that, with SuperPoint feature, our methods perform the best under every error level, even though our models are only trained to classify the 0.1 m, 1 deg error level. When we switch from SuperPoint to SIFT, recall percentages drop for all the methods, however, our data-driven methods still have the highest recall at almost every error level.
To investigate the effectiveness of DINO feature and occlusion handling, we trained several variations of MLP-based and VPT-based scoring models without prefiltering with the occlusion information, as well as VPT-based scoring models without taking DINO features as part of the input. The result shows that, without considering the occlusion, both approaches experience a performance drop, which also occurs in the baseline methods. It is noteworthy that the VPT method can surpass the performance of non-data-driven methods at its trained threshold, even without accounting for occlusion. We accumulate the correctly-localized waypoints along with different distance thresholds and show the result in Figure 6, where it is easier to verify that our learning-based approaches outperform the other baseline methods with both SuperPoint and SIFT features. It is also easy to observe the performance improvement by considering the occlusion.
Results from the 12 waypoints (see Figure 5) in our real-world experiment are shown in Table II. Although the sample size is relatively small, we still see the dominating performance of the VPT-based approach, which shows its generalization ability to the real world. Besides, our approach shows great real-time performance. It achieves to successfully select viewpoints from 100 candidates in less than one second, running on a regular workstation with a NVIDIA GeForce RTX 2080 GPU.
@@
We observe differences between the two learning-based approaches. The transformer based model (VLP) generalises worse between different feature descriptors, but transfers best from simulation data to the real-world tests. The MLP, and the VLP with DINO, perform less good on SuperPoint features and the real-world test, but transfer better to SIFT features. This could indicate that in such situation, they get more semantic prior from DINO, that then is more general accross different features
In conclusion, our proposed data-driven light transformer exhibits optimal performance when evaluated on the same feature descriptors as those used during training. The inclusion of occlusion filtering based on mesh reconstruction enhances the performance of all approaches. It’s important to note that in cases where geometric data isn’t available, the data-driven approach still produces superior results. This could be attributed to the learning process capturing information about the angle ranges from which landmarks were observed, which indirectly encodes geometric information.
@@ VI Conclusion
Future endeavors could focus on further enhancing the generalization of our approach across diverse localization methods, map representations, and environmental conditions. Additionally, there is potential for addressing more complex situations, such as dynamic scenarios, to broaden the scope of applicability.

### v2

@@ IV-B Model Training
We train both two kinds of models on the classification task of predicting whether a viewpoint will result in localization errors smaller than 0.1m and 1 deg. The primary preprocessing actions include normalizing all input features $\mathbf{F}$ to a range between 0 and 1 and ensuring the dataset has a balanced distribution of positive and negative examples. The division of the training and validation datasets is illustrated in Figure 4. For every scene, we randomly generate 100 waypoints. At each waypoint, we sample 50 viewpoints, allowing complete rotation around the yaw axis while keeping the pitch angle within a degree range between -10 to 45. This procedure generates a training and validation dataset comprising 25,000 images and a test set containing 20,000 images.
Figure 4: Overview of the dataset. The number in brackets following the designation corresponds to the index in the Habitat-Matterport 3D dataset. A small collection of scenes lead to great generalization capability of our model, thanks to our effective data point sampling method.
@@ IV-C Evaluation Strategy
As detailed in III, we implement three baseline strategies $\uppi_{\textrm{FIM}},\uppi_{\textrm{max}},\uppi_{\textrm{angle}}$, and augment them with two additional simplistic approaches: $\uppi_{\textrm{forward}}$, which directs the camera towards the next waypoint in the trajectory, and $\uppi_{\textrm{random}}$, which chooses viewpoints at random. Following the methodology established during the training phase, random waypoints and viewpoint samples are generated for each evaluation setting. Each sampled viewpoint is evaluated according to the different viewpoint-selection strategies, recording the localization error for the selected viewpoints. To establish an upper limit for the attainable localization precision at a given position, the minimal error observed from the localization for each sample is also recorded and treated as the pseudo-optimal viewpoint selection strategy.
@@ IV-D Real-world Deployment and Test
Setup To further validate the results obtained from the simulation experiments, the best-performing methods are implemented into a ROS-compatible planning module and deployed on a quadruped robot with a robotic arm that contains a calibrated color camera in its end effector, allowing different viewpoints to be viewed for a given body position. We use a HoloLens 2 to build a map of an indoor environment, as shown in Figure 5.
Data Collection In contrast to the simulated environment where ground-truth poses are readily available, real-world deployments lack these precise positional references. To assess the effectiveness of various viewpoint-choosing strategies in such settings, two kinds of positional data are essential: Each viewpoint planner must first obtain an initial estimate $\hat{\bm{p}}$, which is utilized in 3 to determine an appropriate viewing direction at the current location. Following the selection of a viewing direction, the robot adjusts its arm to capture an image, which is then localized in relation to the pre-existing map. Subsequently, a ground-truth pose is necessary to validate the accuracy of this localization
To measure accurate poses in the map created from the HoloLens recording, we use a combination of visual localization and AprilTag fiducial markers [39]. Once a map of the environment has been created, an AprilTag is affixed to a static location. We capture several high-quality images of the marker, including its surroundings, using a calibrated DSLR camera. The AprilTag marker can be accurately localized within the captured images, and by ensuring that the images contain enough of the surroundings, they can also be localized in the map using visual localization. This two-step process provides an estimated location of the AprilTag marker within the map for each image. An average of the estimated locations is taken in order to minimize the effect of errors in the tag or localization. This process gives us accurate poses of fiducial markers with respect to the captured landmark map.
@@ V Results
Figure 6: Cumulative distribution of position errors of the evaluation points in the testing environments using SuperPoint feature. Methods or versions of methods that rely on the environment mesh in order to determine landmark occlusion are plotted using $\cdot$-$\cdot$-⋅ - ⋅ - lines.
We evaluate our approaches and the baseline methods in both our simulation pipeline and the real world. Table I shows the quantitative results from the simulation. To demonstrate the generalization ability of our approaches across different feature descriptors, we evaluate all methods using both SuperPoint feature [40] and SIFT feature [41], whereas both of our scoring models are only trained with the SuperPoint feature. The result in the table shows that, with SuperPoint feature, our methods perform the best under every error level, even though our models are only trained to classify the 0.1 m, 1 deg error level. When we switch from SuperPoint to SIFT, recall percentages drop for all the methods, however, our data-driven methods still have the highest recall at almost every error level.
To explore the impact of DINO features and occlusion handling on model performance, we conduct experiments with various model configurations, including MLP-based and VPT-based models that do not pre-filter based on occlusion information, as well as VPT-based models that do not incorporate DINO features in their input. The findings reveal that neglecting occlusion leads to diminished performance for both model types, a trend that is also observed in baseline methods. Interestingly, the VPT-based model demonstrates the ability to exceed the efficacy of traditional non-data-driven approaches at its designated training threshold, even without occlusion considerations. We accumulate the correctly-localized waypoints along with different distance thresholds and show the result in Figure 6, where it is easier to verify that our learning-based approaches outperform the other baseline methods with both SuperPoint and SIFT features.
Results from the 12 waypoints (see Figure 5) in our real-world experiment are shown in Table II. Although the sample size is relatively small, we still see the dominating performance of the VPT-based model, which shows its generalization ability to the real world. Besides, our final approach shows great online performance. It achieves to successfully select viewpoints from 100 candidates in less than one second, running on a regular workstation with a NVIDIA GeForce RTX 2080 GPU.
@@
We note variations in performance between the two machine learning methodologies. The transformer model exhibits reduced generalization across distinct feature descriptors yet demonstrates superior transferability from simulated data to real-world applications. Conversely, the MLP-based model and the VLP-based model employing DINO underperform with SuperPoint features and in real-world evaluations but show enhanced adaptability to SIFT features. This suggests that in these scenarios, they benefit from a more substantial semantic prior provided by DINO, which in turn offers greater generality across various feature types.
In conclusion, our proposed data-driven light transformer-based model exhibits optimal performance when evaluated on the same feature descriptors as those used during training. The inclusion of occlusion filtering based on mesh reconstruction enhances the performance of all approaches. It’s important to note that in cases where geometric data isn’t available, the proposed data-driven approach still produces superior results. This could be attributed to the learning process capturing information about the angle ranges from which landmarks were observed, which indirectly encodes geometric information.